OpenStudy (anonymous):

f =1/(2pi squareroot (L*C)) Solve for L

OpenStudy (anonymous):

I got L = 1/C*4pi^2*f^2 is this correct?

OpenStudy (anonymous):

\[f _{0}=\frac{ 1 }{ 2\pi \sqrt{LC} }\]

OpenStudy (anonymous):

^ original question

OpenStudy (anonymous):

\[L=\frac{ 1 }{ C4\pi^2f _{0}^2 }\] My answer

OpenStudy (anonymous):

Yes. You did correctly

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Looks good to me.